OpenStudy (anonymous):

In covering a distance of 30 km, Arun takes 2 hours more than Anil. If Arun doubles his speed, then he would take 1 hour less than Anil. What is Arun's speed?

OpenStudy (anonymous):

dist=speed*time so time =x speed=y then 1---> (x+2)y=30 2--->(x-1)2y=30

OpenStudy (anonymous):

yx=10,y=10,x=1

OpenStudy (anonymous):

divide 1 and 2 then solve for x first...

OpenStudy (anonymous):

now use x= 3 in any of the equation to get y =6...

OpenStudy (anonymous):

no x=-4

OpenStudy (anonymous):

y=-15

OpenStudy (anonymous):

check it @matricked

OpenStudy (anonymous):

yup my calculation is wrong ...

OpenStudy (anonymous):

x=4 dear....

OpenStudy (anonymous):

(x+2)=2(x-1) x+2=2x -2 x-2x=-2-2 -x=-4 hence x=4

OpenStudy (anonymous):

y= 5

OpenStudy (anonymous):

thanks for correcting my calculation...

OpenStudy (anonymous):

you are welcome!

OpenStudy (anonymous):

plse tell me matrick why i get different answer for this eqn 1 yx+2y=30 2yx-2y=30 so yx=10,y=10 if we substitute in eqn 1and y=5 in eqn 2

OpenStudy (anonymous):

if xy =10 then y=10 and x=1 are not the only factors...of 10 and not only that x and y may be integers...

OpenStudy (anonymous):

in case if xy =0 then only we can xay either x=0 or y=0 or both..

OpenStudy (anonymous):

so by using ur method we are getting the right answer.but if y=5,x=4 this is not equal to 10 it is equal to 20

OpenStudy (anonymous):

something went wrong in eqn 1

OpenStudy (anonymous):

its really confusing .but i got it by different method